OpenStudy (ksaimouli):

a building is 10.0m tall.From the top of the building a stone is dropped from rest.when the stone is halfway down a second stone is thrown straight downward from the top of the building.the two stones strike the ground simultaneously. find the initial speed of the second stone

OpenStudy (anonymous):

Isn't this rectilinear motion? Physics?

OpenStudy (ksaimouli):

kinematics

OpenStudy (anonymous):

good meters g= -9.8m/s x_o=10 first solve for t when the first stone reaches 5 m

OpenStudy (anonymous):

\[h=-\frac{g t^2}{2}-vt\]

OpenStudy (anonymous):

\[+h_0\]

OpenStudy (ksaimouli):

what i did is i found the time for first stone and divided by 2

OpenStudy (ksaimouli):

so i will get T for second stone

OpenStudy (anonymous):

cant, acceleration

OpenStudy (anonymous):

first determine the time when the first rock reaches half the height of the building so 5=(1/2)(-9.8)t^2+10 then determine when the first stone reaches the ground 0=(1/2)(-9.8)t^2+10 take the difference of the 2 times thats how long it takes the 2nd stone to reach the ground

OpenStudy (anonymous):

so you then solve for v_o of the 2nd stone 0=(1/2)(-9.8)t^2+v_o*t+10 where t is the difference in time of the first stone

OpenStudy (ksaimouli):

but a=9.8 right not (-)

OpenStudy (anonymous):

i use g=-9.8 because its moving in a downwards motion

OpenStudy (anonymous):

it doesnt matter if its positive or negative, as long as the directions are the same it would be -9.8 because we start at a height, as time passes, the height becomes smaller, thus acceleration must be negative, to reduce the height

OpenStudy (anonymous):

it would be positive if you used -10 instead of +10 if you used a postive accelereration, then the initial velocity would have to be positive because they are both travelling in the same direction

OpenStudy (ksaimouli):

i got t=1.01s for the first stone

OpenStudy (ksaimouli):